# Cutting Through the Noise: Boosting LLM Performance on Math Word Problems

### v3

@@ Abstract
We introduce a dataset, ProbleMathic, containing both adversarial and non-adversarial MWPs.
@@
To mitigate this, we fine-tune LLMs (Llama-2, Mistral) on the adversarial samples from our dataset.
@@
Dataset: https://huggingface.co/datasets/him1411/problemathic
@@ 1 Introduction
Large Language Models (LLMs) AI@Meta (2024); Team et al. (2023); Achiam et al. (2023) have achieved impressive performance on various tasks, including mathematical problem solving Imani et al. (2023); Gaur and Saunshi (2023); Romera-Paredes et al. (2023); Ahn et al. (2024). Existing mathematical datasets Patel et al. (2021); Sawada et al. (2023); Mishra et al. (2022); Gupta et al. (2023) typically feature simplified questions with limited variables and numerical data directly relevant to the problem. However, real-world MWPs arise within the context of broader reasoning tasks. Thus, they are likely to contain irrelevant information stemming from the wider text. This unrelated information acts as noise, distracting the reasoning process of language models (Fig. 1). Recent works have shown that LLMs do not perform consistently well on mathematical reasoning Ahn et al. (2024); Shakarian et al. (2023); Chang et al. (2023) and are sensitive to linguistic variations in MWPs Patel et al. (2021); Shi et al. (2023); Kumar et al. (2021). Previous works Kumar et al. (2022) have attempted to make LLMs more robust by generating "adversarial" training data but fail to emulate real-world data by not introducing textual noise.
@@
| Mistral Large | 92.50 | 90.74 | 69.72 | 92.50 | 75.40 | 89.81 | 65.13 | 93.75 | 73.77 | 81.48 | 70.64 | |
| Claude-3 Sonnet | 57.37 | 87.96 | 68.80 | 93.75 | 68.85 | 86.11 | 75.22 | 96.25 | 67.21 | 82.40 | ||
| Reka Flash (21B) | 62.29 | 79.63 | 64.22 | 95.00 | 60.65 | 83.33 | 55.05 | 93.75 | 65.57 | 84.26 | 63.30 | |
@@
To initiate a systematic study, we introduce a new dataset ProbleMathic containing both adversarial and non-adversarial mathematical word problem (MWP) pairs.
The dataset includes problems of two difficulty levels: Simple and Complex.
These adversarial MWPs are augmented to add irrelevant numerical information while maintaining the integrity of the original question; i.e. their solution and final answer are the same as their non-adversarial counterparts.
We employ various LLMs, including Gemini-1.5 Pro, Claude-3 Sonnet, and Qwen-1.5, among others, for inference.
Our observations reveal significant performance drops (16.07%, 22.93%, 41.53%, respectively) between adversarial and non-adversarial problems, highlighting the models’ inability to reason accurately with noisy data.
To address this challenge, we propose a prompting framework that allows us to leverage existing MWPs of varying complexities and generate augmented adversarial versions for the training set.
We demonstrate that fine-tuning LLMs on these samples enhances their ability to identify irrelevant variables.
Fine-tuning Llama-2 (7B, 13B) and Mistral (7B) on these samples yields an average 8% improvement in performance on both Simple (8.19%) and Complex (7.07%) adversarial test samples.
Finally, to assess the generalizability of our prompting framework, we create an adversarial variant of the widely used GSM-8K benchmark (GSM-8K-Adv).
This reveals that LLMs (Mistral Large, Llama-2, and Llama-3) experience an average performance drop of when tested on this adversarial benchmark.
Additionally, we compare our approach with others in the domain of adversarial data generation in §C.
Contributions:
(a) We introduce ProbleMathic, a dataset of Simple and Complex MWPs that demonstrates the susceptibility of LLMs to irrelevant numerical information. (b) We propose a prompting framework to generate adversarial variants of existing MWPs. We show that fine-tuning on these adversarial samples leads to improved performance and robustness to noise.
(c) Using this prompting framework, we introduce GSM-8K-Adv, an adversarial variant of the GSM-8K benchmark.
2 Methodology
2.1 Adversarial data generation
We propose a constrained additive approach, wherein we add numerical information (Structural Invariance) to the original MWP. We use our hand-crafted set of MWPs as seed samples and augment them in accordance with the following constraints: (1) The added variables must not be related to or derived from the existing variables in the passage. (2) The added variables must not share the same physical unit as any of the original variables. (3) The augmented text should not add any numerical information about existing variables that did not exist in the original passage. This gives us a set of adversarial MWPs containing variables that do not feature in their solution equations. Hence, this allows us to evaluate the model’s ability to identify relevant information and demonstrate reasoning ability beyond logically connecting all numerical entities in the input in order to produce a solution. The prompting framework used to adversarially augment the seed samples can be found in §F. We also compare our adversarial data generation with other contemporaneous approaches (§C).
2.2 Dataset
ProbleMathic is split into Simple and Complex problem sets, containing both adversarial and non-adversarial samples. The Simple set consists of MWPs requiring a single mathematical operation to solve. Number of samples in train and test splits are 1313 and 60 respectively. These samples are aggregated from existing sources detailed in §D. The Complex set contains MWPs involving more than one mathematical operation. Number of samples in train and test splits are 613 and 73 respectively. Table 1 highlights some examples of train and test set of the dataset. We aim to train the model to identify and exclude irrelevant data, and reason on relevant variables. To this end, we use GPT-4 Achiam et al. (2023) to generate solution explanations for all training samples. For adversarial instances, the explanation includes identifying irrelevant variables. The adversarial test instances are hand-crafted to eschew any data leakage (§D.1).
@@ 2.2 Dataset
| Llama-3 (70B) | 86.59 | 18.41 | |
| Mistral Large | 21.38 | ||
| Claude-3 Sonnet | 69.39 | 22.93 | |
@@
| Zero-shot | One-shot | Two-shot | ||||
| Og | Adv | Og | Adv | Og | Adv | |
| Yi (34B) | 42.21 | 38.13 | 54.31 | 38.33 | 53.34 | 50.87 |
| Llama-2 (70B) | 38.54 | 36.17 | 32.33 | 29.97 | 34.58 | 30.28 |
| Llama-3 (70B) | 65.36 | 60.01 | 69.29 | 66.31 | 72.64 | 68.06 |
| Mistral Large | 71.38 | 65.64 | 74.14 | 69.86 | 78.93 | 74.68 |
| Qwen-1.5 (72B) | 59.23 | 53.56 | 65.94 | 62.04 | 68.21 | 63.32 |
3 Experiments and Results
3.1 Experimental setup
We conduct zero-shot, one-shot and two-shot inference experiments with Claude-3 Sonnet Anthropic (2024), Gemini-1.5 Pro Team et al. (2023), Reka Flash Ormazabal et al. (2024), Yi (34B) AI et al. (2024), Llama-2 (70B) Touvron et al. (2023), Llama-3 (70B) AI@Meta (2024), Mistral Large AI (2024), and Qwen-1.5 (72B) Bai et al. (2023) to highlight these models’ sensitivity to adversarial noise. We then conduct fine-tuning experiments on Llama-2 (7B, 13B) Touvron et al. (2023) and Mistral (7B) Jiang et al. (2023). These models are fine-tuned in 3 settings for Simple and Complex problems each: (1) fine-tuning using non-adversarial samples, (2) using adversarial samples, and (3) combined training where we fine-tune over a combined set of non-adversarial and adversarial samples. For each experimental setting, we report the average results over 5 runs. We use exact match accuracy as our evaluation metric.
3.2 Results
3.2.1 Zero-shot and few-shot inference
@@ 3.2.2 Fine-tuning
Table 4 presents LLM fine-tuning results across various training data scenarios: Og (non-adversarial), adversarial, and a combined set. Performance is evaluated on Og and adversarial samples from both Simple and Complex problem sets. Llama-2 (13B) is the best performer, likely due to its large size. The optimal fine-tuning setting, highlighted in varies across test sets. While similar samples yield the best performance on Og Simple problems, training on adversarial samples boosts average performance on adversarial Simple problems and both Og and adversarial Complex problems by 8.19%, 7.07%, and 1.78%, respectively. Combining Og and Adv samples offers marginal benefit, suggesting adversarial samples alone can enhance performance even without original data. We also present fine-tuning results with these models on the MultiArith Roy and Roth (2016) dataset in §B.
4 GSM-8K-Adv: an adversarial variant of an existing benchmark
To show the generalizability of our prompting strategy (§2.1) to create adversarial test sets, we use it to create GSM-8K-Adv, an adversarial variant of GSM-8K dataset. We test models on the original benchmark as well as our adversarial variant under zero-shot, one-shot, and two-shot settings. The results in Table 5 echo our observations on the ProbleMathic dataset: models struggle when faced with adversarial information, resulting in a performance drop of .
5 Conclusion
Limitations
Our study presents a dataset comprising approximately 1500 MWPS divided into two difficulty levels. Additionally, we introduce an adversarial variant of the dataset featuring irrelevant variables, along with a framework for generating these adversarial samples. Our fine-tuning efforts were limited to 13B models to observe incremental improvements; fine-tuning larger models (>40B) was not feasible due to resource limitations. While we have introduced a substantially challenging dataset, creating a dataset with higher difficulty levels and more adversaries could further challenge LLMs. Future investigations could delve into generating and incorporating more nuanced perturbations in MWPs to heighten the challenge of adversarial problem-solving. Furthermore, resource constraints prevented the reproduction of adversarial versions of other popular benchmarks beyond the GSM-8K dataset in our study. A promising avenue for future research could involve devising challenging adversarial variants of other widely-used benchmarks.
Appendix A Hyperparameters
Appendix B Results on MultiArith dataset
| Test set | MultiArith (Og) | MultiArith (Adv) | ||
| Training set | Og | Adv | Og | Adv |
| Mistral (7B) | 29.21 | 30.03 | 27.47 | 35.87 |
| Llama-2 (7B) | 24.56 | 26.43 | 10.32 | 23.91 |
| Llama-2 (13B) | 36.75 | 32.21 | 30.68 | 37.21 |
Appendix C Comparison with other adversarial generation frameworks
| Adversarial Examples for Evaluating Reading Comprehension Systems Jia and Liang (2017) | Text | SQuAD | Rsn. | Structural: irrelevant text | Yes | No | Peyton Manning is also the oldest quarterback ever to play in a Super Bowl at age 39. John Elway, who led the Broncos to victory in Super Bowl XXXIII at age 38, is currently Denver’s Executive VP of Football Operations. Quarterback Jeff Dean had jersey 37 in Champ Bowl XXXIV. Which quarterback was 38 in Super Bowl XXXIII? |
We compare ProbleMathic with other adversarial generation mthods along various facets, including their generated adversarial variant on the sample MWP shown in Fig 1. The axes of comparison we show are:
@@ Appendix C Comparison with other adversarial generation frameworks
C.1 Comparison with "Adversarial Math Word Problem Generation"
@@
| Model | GSM8K | GSM8K* | GSM8K-Adv |
| Yi (34B) | 42.21 | 41.93 | 38.13 |
| Llama 2 (70B) | 38.54 | 38.72 | 36.17 |
| Llama 3 (70B) | 65.36 | 65.4 | 60.01 |
| Mistral Large | 71.38 | 70.12 | 65.64 |
| Qwen-1.5 (72B) | 59.23 | 60.01 | 53.56 |
@@
A work that has drawn comparisons to ProbleMathic is that of Xie et al. (2024). Our framework differ from this approach in various salient aspects. Xie et al. (2024) change the values present in the original problem without altering the structure or number of variables, posing little challenge to self-consistent models. Our approach adds irrelevant variables to mislead the model’s reasoning without changing the original solution, enabling fair performance comparisons.The adversaries generated by Xie et al. (2024) replace the original values with substituted values, testing LLMs’ memorization abilities and showing that models falter when faced with identical problems with changed values. However, sufficiently powerful and self-consistent models can reason over similar problems and remain coherent (Table 8). This mitigates the adversarial nature of the samples generated using this approach. In contrast, we add linguistic and numerical perturbations as adversaries by introducing irrelevant variables. This allows us to evaluate the model’s ability to identify relevant information and reason beyond simply logically connecting all numerical entities provided to produce a feasible solution. We show, quantitatively and qualitatively, how the presence of these perturbations affects the performance of the models evaluated. Finally, Xie et al. (2024) ASR (Attack Success Rate) to measure adversarial efficacy, showing 1 in 100 numerically-perturbed variants results in an incorrect solution. Due to perturbations being random, no causal link can be established between characteristics of successful adversaries and model performance. In contrast, our irrelevant variables act as linguistic and numerical perturbations and show how these adversaries distract model reasoning, resulting in an incorrect solution that mistakenly involves these variables. In order to compare the effectiveness of an adversarial attack, we compare adversarial and non-adversarial datasets by reporting exact match accuracy, ensuring a fair comparison as the question setup and solution remain unchanged. We generate an adversarial variant per Xie et al. (2024) for the MWP shown in Fig 1 of our paper and show LLM outcomes on all samples.
@@
To systematically show that the samples in ProbleMathic are more adversarial, we compare exact match accuracy of zeroshot inference on GSM8K (original), GSM8K-Adv (our variant), and the adversarial GSM8K shared by Xie et al. (2024) we denote as GSM8K* (Table 9). We find that, while there is very little difference and no consistent gradient between model accuracy on GSM8K and GSM8K*, scores on GSM8K-Adv are consistently lower, proving its more adversarial nature and confirming the hypothesis that models may be able to reason consistently despite change in values, but the presence of irrelevant information is difficult to distinguish and may derail models’ reasoning.
Appendix D Curating ProbleMathic
Building upon prior efforts in mathematical problem dataset curation, we leverages existing MWP datasets Koncel-Kedziorski et al. (2016); Roy and Roth (2016, 2018); Kushman et al. (2014). These MWPs are aggregated, undergo data cleaning procedures, and are subsequently categorized into distinct Simple and Complex subsets based on the number and complexity of mathematical operations required to reach the solution.
D.1 Human-crafted test samples
In the human-crafted test set, we utilize various categories of adversaries. These categories and their frequency of occurrence are listed in Table 10.
| Type | Sample | Explanation | Samples (Simp.) | Samples (Compl.) |
Appendix E Human quality evaluation
Appendix F Prompting Framework for Adversarial Data Generation

### v5

@@ Abstract
We introduce a dataset, ProbleMathic containing both adversarial and non-adversarial MWPs.
@@
To mitigate this, we fine-tune LLMs (Qwen-2, Mistral) on the adversarial samples from our dataset.
@@
Dataset: https://huggingface.co/datasets/him1411/problemathic .
@@ 1 Introduction
Large Language Models (LLMs) (AI@Meta, 2024; Team et al., 2023; Achiam et al., 2023) have achieved impressive performance on various tasks, including mathematical problem solving (Imani et al., 2023; Gaur & Saunshi, 2023; Romera-Paredes et al., 2023; Ahn et al., 2024). Existing mathematical datasets (Patel et al., 2021; Sawada et al., 2023; Mishra et al., 2022; Gupta et al., 2023) typically feature simplified questions with limited variables and numerical data directly relevant to the problem. However, real-world MWPs arise within the context of broader reasoning tasks. Thus, they are likely to contain irrelevant information stemming from the wider text. This unrelated information acts as noise, distracting the reasoning process of language models (Fig. 1). Recent works have shown that LLMs do not perform consistently well on mathematical reasoning (Ahn et al., 2024; Shakarian et al., 2023; Chang et al., 2023) and are sensitive to linguistic variations in MWPs (Patel et al., 2021; Shi et al., 2023; Kumar et al., 2021). Previous works (Kumar et al., 2022) have attempted to make LLMs more robust by generating "adversarial" training data but fail to emulate real-world data by not introducing textual noise.
To initiate a systematic study, we introduce a new dataset ProbleMathic containing both adversarial and non-adversarial mathematical word problem (MWP) pairs. The dataset includes problems of two difficulty levels: Simple and Complex. These adversarial MWPs are augmented to add irrelevant numerical information while maintaining the integrity of the original question; i.e. their solution and final answer are the same as their non-adversarial counterparts. We employ various LLMs, including Gemini-1.5 Pro, Claude-3.5 Sonnet, and Qwen-1.5, among others, for inference. Our observations reveal significant performance drops (16.07%, 22.93%, 41.53%, respectively) between adversarial and non-adversarial problems, highlighting the models’ inability to reason accurately with noisy data. To address this challenge, we propose a prompting framework that allows us to leverage existing MWPs of varying complexities and generate augmented adversarial versions for the training set. We demonstrate that fine-tuning LLMs on these samples enhances their ability to identify irrelevant variables. Fine-tuning Llama-2 (7B, 13B) and Mistral (7B) on these samples yields an average 8% improvement in performance on both Simple (8.19%) and Complex (7.07%) adversarial test samples. Finally, to assess the generalizability of our prompting framework, we create an adversarial variant of the widely used GSM-8K benchmark (GSM-8K-Adv). This reveals that LLMs (Mistral Large, Llama-2, and Llama-3) experience an average performance drop of when tested on this adversarial benchmark. Additionally, we compare our approach with others in the domain of adversarial data generation in §4.
Contributions: (a) We introduce ProbleMathic, a dataset of Simple and Complex MWPs that demonstrates the susceptibility of LLMs to irrelevant numerical information. (b) We propose a prompting framework to generate adversarial variants of existing MWPs. We show that fine-tuning on these adversarial samples leads to improved performance and robustness to noise. (c) Using this prompting framework, we introduce GSM-8K-Adv, an adversarial variant of the GSM-8K benchmark.
2 Methodology
2.1 Adversarial data generation
We propose a constrained additive approach, wherein we add numerical information (Structural Invariance) to the original MWP. We use our hand-crafted set of MWPs as seed samples and augment them in accordance with the following constraints: (1) The added variables must not be related to or derived from the existing variables in the passage. (2) The added variables must not share the same physical unit as any of the original variables. (3) The augmented text should not add any numerical information about existing variables that did not exist in the original passage. This gives us a set of adversarial MWPs containing variables that do not feature in their solution equations. Hence, this allows us to evaluate the model’s ability to identify relevant information and demonstrate reasoning ability beyond logically connecting all numerical entities in the input in order to produce a solution. The prompting framework used to adversarially augment the seed samples can be found in §C. We also compare our adversarial data generation with other contemporaneous approaches (§4).
| Sample Type | Sample | |||||||||||||
2.2 Dataset
| Type | Sample | Explanation | Samples (Simp.) | Samples (Compl.) |
ProbleMathic is split into Simple and Complex problem sets, containing both adversarial and non-adversarial samples. Building upon prior efforts in mathematical problem dataset curation, we leverage existing MWP datasets (Koncel-Kedziorski et al., 2016; Roy & Roth, 2016; 2018; Kushman et al., 2014). These MWPs are aggregated, undergo data cleaning procedures, and are subsequently categorized into distinct Simple and Complex subsets based on the number and complexity of mathematical operations required to reach the solution. We distinguish between Simple and Complex problems as follows:
Simple: MWPs involving no more than 3 linear operations. These problems require addition and subtraction operations only. Number of samples in train and test splits are 1313 and 60 respectively.
Complex: MWPs involving multiple nested operations. These questions often involve rates and proportions. They require multiplication and division operations too. Number of samples in train and test splits are 613 and 73 respectively.
Table 1 highlights some examples of train and test set of the dataset. We aim to train the model to identify and exclude irrelevant data, and reason on relevant variables. To this end, we use GPT-4 (Achiam et al., 2023) to generate solution explanations for all training samples. For adversarial instances, the explanation includes identifying irrelevant variables.
The adversarial test instances are hand-crafted to eschew any data leakage. In the human-crafted test set, we utilize various categories of adversaries. These categories and their frequency of occurrence are listed in Table 2.
To show the generalizability of our prompting strategy (§2.1) to create adversarial test sets, we use it to create GSM-8K-Adv, an adversarial variant of GSM-8K dataset.
2.3 Human quality evaluation
3 Experiments and Results
3.1 Experimental setup
We conduct zero-shot, one-shot and two-shot inference experiments with Claude-3 Sonnet (Anthropic, 2024), Gemini-1.5 Pro (Team et al., 2023), Reka Flash (Ormazabal et al., 2024), Yi (34B) (AI et al., 2024), Llama-2 (70B) (Touvron et al., 2023), Llama-3 (70B) (AI@Meta, 2024), Mistral Large (AI, 2024), and Qwen-1.5 (72B) (Bai et al., 2023) to highlight these models’ sensitivity to adversarial noise. We then conduct fine-tuning experiments on Llama-2 (7B, 13B) (Touvron et al., 2023) and Mistral (7B) (Jiang et al., 2023). These models are fine-tuned in 3 settings for Simple and Complex problems each: (1) fine-tuning using non-adversarial samples, (2) using adversarial samples, and (3) combined training where we fine-tune over a combined set of non-adversarial and adversarial samples. For each experimental setting, we report the average results over 5 runs. We use exact match accuracy as our evaluation metric.
3.2 Results
3.2.1 Zero-shot and few-shot inference
@@
| Mistral Large | 92.50 | 70.49 | 90.74 | 69.72 | 92.50 | 75.40 | 89.81 | 65.13 | 93.75 | 73.77 | 81.48 | 70.64 |
| Claude-3 Sonnet | 93.75 | 57.37 | 87.96 | 68.80 | 93.75 | 68.85 | 86.11 | 75.22 | 96.25 | 67.21 | 82.40 | 78.89 |
| Claude-3.5 Sonnet | 64.87 | 88.45 | 74.65 | 95.26 | 72.65 | 88.44 | 78.56 | 71.23 | 85.15 | |||
| Llama-3.3 (70B) | 89.38 | 79.95 | 64.63 | 94.70 | 75.97 | 83.68 | 70.40 | 97.09 | 79.79 | 84.21 | 72.34 | |
@@
| Llama-3 (70B) | 86.59 | 70.65 | 18.41 |
| Mistral Large | 90.13 | 70.86 | 21.38 |
| Claude-3 Sonnet | 90.04 | 69.39 | 22.93 |
| Claude-3.5 Sonnet | 91.72 | 73.54 | 19.82 |
| Llama-3.3 (70B) | 88.16 | 72.44 | 17.83 |
| Llama-2 (70B) | 31.02 | ||
@@
| Zero-shot | One-shot | Two-shot | ||||
| Og | Adv | Og | Adv | Og | Adv | |
| Claude-3.5 Sonnet | 93.6* | 70.81 | 94.81 | 80.32 | 95.32 | 83.19 |
| Claude-3.5 Haiku | 88.9* | 64.33 | 90.12 | 72.55 | 92.01 | 75.06 |
| Yi (34B) | 42.21 | 38.13 | 54.31 | 38.33 | 53.34 | 50.87 |
| Llama-2 (70B) | 38.54 | 36.17 | 32.33 | 29.97 | 34.58 | 30.28 |
| Llama-3 (70B) | 65.36 | 60.01 | 69.29 | 66.31 | 72.64 | 68.06 |
| Llama-3.3 (70B) | 90.12 | 71.42 | 92.70 | 75.39 | 92.19 | 77.19 |
| Mistral Large | 71.38 | 65.64 | 74.14 | 69.86 | 78.93 | 74.68 |
| Qwen-1.5 (72B) | 59.23 | 53.56 | 65.94 | 62.04 | 68.21 | 63.32 |
| Qwen2-MI (72B) | 96.7* | 73.42 | 96.92 | 83.37 | 96.45 | 84.02 |
@@
| Qwen-2.5 (14B) | 85.21 | 87.32 | 80.43 | 83.82 | 72.56 | 72.76 | 70.18 | 71.43 | ||||
| Phi-4 (14B) | 84.37 | 87.51 | 84.26 | 85.94 | 80.34 | 81.43 | 84.52 | 85.02 | ||||
@@ 3.2.2 Fine-tuning
Table 6 presents LLM fine-tuning results across various training data scenarios: Og (non-adversarial), adversarial, and a combined set. Performance is evaluated on Og and adversarial samples from both Simple and Complex problem sets. Llama-2 (13B) is the best performer, likely due to its large size. The optimal fine-tuning setting, highlighted in varies across test sets. While similar samples yield the best performance on Og Simple problems, training on adversarial samples boosts average performance on adversarial Simple problems and both Og and adversarial Complex problems by 8.19%, 7.07%, and 1.78%, respectively. Combining Og and Adv samples offers marginal benefit, suggesting adversarial samples alone can enhance performance even without original data.
3.3 Adversarial variants of existing datasets
Additionally, we test models on the original GSM8K benchmark as well as our adversarial variant (GSM8K-Adv) under zero-shot, one-shot, and two-shot settings. The results in Table 5 echo our observations on the ProbleMathic dataset: models struggle when faced with adversarial information, resulting in a performance drop of .
4 Related Work
We compare ProbleMathic with other adversarial generation methods along various facets, including their generated adversarial variant on the sample MWP shown in Fig 1. The axes of comparison we show are:
@@ 4 Related Work
| Adversarial Examples for Evaluating Reading Comprehension Systems Jia & Liang (2017) | Text | SQuAD | Rsn. | Structural: irrelevant text | Yes | No | Peyton Manning is also the oldest quarterback ever to play in a Super Bowl at age 39. John Elway, who led the Broncos to victory in Super Bowl XXXIII at age 38, is currently Denver’s Executive VP of Football Operations. Quarterback Jeff Dean had jersey 37 in Champ Bowl XXXIV. Which quarterback was 38 in Super Bowl XXXIII? |
4.1 Comparison with "Adversarial Math Word Problem Generation"
@@
A work that has drawn comparisons to ProbleMathic is that of Xie et al. (2024). Our framework differs from this approach in various salient aspects. Xie et al. (2024) change the values present in the original problem without altering the structure or number of variables, posing little challenge to self-consistent models. Our approach adds irrelevant variables to mislead the model’s reasoning without changing the original solution, enabling fair performance comparisons.The adversaries generated by Xie et al. (2024) replace the original values with substituted values, testing LLMs’ memorization abilities and showing that models falter when faced with identical problems with changed values. However, sufficiently powerful and self-consistent models can reason over similar problems and remain coherent (Table 8). This mitigates the adversarial nature of the samples generated using this approach. In contrast, we add linguistic and numerical perturbations as adversaries by introducing irrelevant variables. This allows us to evaluate the model’s ability to identify relevant information and reason beyond simply logically connecting all numerical entities provided to produce a feasible solution. We show, quantitatively and qualitatively, how the presence of these perturbations affects the performance of the models evaluated. Finally, Xie et al. (2024) ASR (Attack Success Rate) to measure adversarial efficacy, showing 1 in 100 numerically-perturbed variants results in an incorrect solution. Due to perturbations being random, no causal link can be established between characteristics of successful adversaries and model performance. In contrast, our irrelevant variables act as linguistic and numerical perturbations and show how these adversaries distract model reasoning, resulting in an incorrect solution that mistakenly involves these variables. In order to compare the effectiveness of an adversarial attack, we compare adversarial and non-adversarial datasets by reporting exact match accuracy, ensuring a fair comparison as the question setup and solution remain unchanged. We generate an adversarial variant per Xie et al. (2024) for the MWP shown in Fig 1 of our paper and show LLM outcomes on all samples.
@@
To systematically show that the samples in ProbleMathic are more adversarial, we compare exact match accuracy of zeroshot inference on GSM8K (original), GSM8K-Adv (our variant), and the adversarial GSM8K shared by (Xie et al., 2024) we denote as GSM8K* (Table 5). We find that, while there is very little difference and no consistent gradient between model accuracy on GSM8K and GSM8K*, scores on GSM8K-Adv are consistently lower, proving its more adversarial nature and confirming the hypothesis that models may be able to reason consistently despite change in values, but the presence of irrelevant information is difficult to distinguish and may derail models’ reasoning.
5 Conclusion
Appendix
Appendix A Hyperparameters
Appendix B Results on MultiArith dataset
We also present fine-tuning results with these models on the MultiArith (Roy & Roth, 2016) dataset.
| Test set | MultiArith (Og) | MultiArith (Adv) | ||
| Training set | Og | Adv | Og | Adv |
| Mistral (7B) | 29.21 | 30.03 | 27.47 | 35.87 |
| Llama-2 (7B) | 24.56 | 26.43 | 10.32 | 23.91 |
| Llama-2 (13B) | 36.75 | 32.21 | 30.68 | 37.21 |
We denote the adversarial and non-adversarial variants by Adv and Og respectively. The results are shown in Table 9.
Appendix C Prompting Framework for Adversarial Data Generation